# FastVID: Dynamic Density Pruning for Fast Video Large Language Models

### v1

@@ 4 Experiments
Model Method Retention Ratio $R$ MVBench LongVideo Bench MLVU VideoMME Avg.
@@
LLaVA- OneVision [14] Vanilla 100% 56.9 56.4 65.2 58.6 70.3 56.6 48.8 59.3 100
DyCoke [31] 32% 56.3 56.6 62.1 57.1 68.1 56.7 46.7 58.0 97.8
DyCoke [31] 25% 49.5 48.1 55.8 51.0 61.1 48.6 43.2 51.1 86.2
FastV [4] 54.7 55.5 61.5 56.2 68.0 54.6 46.0 57.0 96.1
VisionZip [38] 53.7 51.2 58.5 54.1 61.6 53.4 47.2 54.4 91.7
VisionZip [38] 56.6 55.7 64.8 58.0 68.6 57.7 47.7 58.8 99.1
PruneVID [11] 55.0 55.4 64.6 57.5 69.7 54.8 48.1 58.1 98.0
FastVID 56.5 56.3 64.1 58.0 69.9 56.6 47.7 58.7 99.0
FastV [4] 20% 54.1 56.6 61.2 56.2 66.8 54.6 47.2 57.0 96.1
VisionZip [38] 53.0 50.0 57.1 53.0 60.8 51.0 47.1 53.3 90.0
VisionZip [38] 55.8 55.4 64.2 58.0 68.6 57.0 48.3 58.4 98.5
PruneVID [11] 54.6 54.7 63.7 56.2 68.1 53.9 46.7 57.3 96.6
FastVID 56.3 57.1 63.9 57.9 69.3 56.7 47.7 58.8 99.1
FastV [4] 15% 53.2 54.9 59.8 54.7 65.1 53.4 45.7 55.7 93.9
VisionZip [38] 50.3 46.9 54.4 49.5 55.8 49.3 43.3 50.3 84.8
VisionZip [38] 54.3 53.9 63.1 55.5 63.0 54.4 49.1 56.7 95.6
PruneVID [11] 54.6 53.6 63.3 56.1 67.6 53.4 47.2 56.9 96.0
FastVID 56.0 56.2 63.2 57.7 69.3 56.2 47.4 58.3 98.3
FastV [4] 10% 51.7 52.1 57.7 52.4 60.9 51.4 45.0 53.5 90.2
VisionZip [38] 44.4 43.5 51.5 46.0 50.4 45.8 41.8 46.4 78.3
VisionZip [38] 51.7 48.3 59.7 52.8 59.4 52.0 46.9 53.1 89.6
PruneVID [11] 53.2 54.4 61.7 56.0 66.6 53.0 48.4 56.3 94.9
FastVID 55.9 56.3 62.7 57.3 67.4 56.0 48.6 58.1 98.0
LLaVA- Video [43] Vanilla 100% 60.4 59.6 70.3 64.1 76.9 62.1 53.4 63.6 100
DyCoke [31] 32% 59.3 57.9 65.7 61.6 74.6 59.0 51.1 61.1 96.1
DyCoke [31] 25% 50.8 53.0 56.9 56.1 65.8 53.6 48.9 54.2 85.2
FastV [4] 58.0 58.3 63.9 61.0 71.3 60.7 51.0 60.3 94.8
VisionZip [38] 56.4 54.1 62.1 58.6 66.3 58.2 51.2 57.8 90.9
VisionZip [38] 58.3 58.3 66.6 61.9 73.3 60.1 52.2 61.5 96.7
PruneVID [11] 55.3 57.7 64.1 60.2 72.4 57.7 50.6 59.3 93.2
FastVID 59.9 57.4 68.6 63.6 74.9 62.2 53.7 62.4 98.1
FastV [4] 10% 55.8 55.4 58.9 57.9 67.6 57.7 48.6 57.0 89.6
VisionZip [38] 46.3 46.6 52.2 49.5 54.2 49.9 44.3 48.7 76.6
VisionZip [38] 56.6 53.6 61.7 58.7 67.6 58.6 50.1 57.7 90.7
PruneVID [11] 54.3 55.3 61.8 56.4 67.9 54.1 47.1 57.0 89.6
FastVID 58.5 56.5 64.9 60.7 71.7 59.1 51.2 60.2 94.7
Table 1: Comparison of state-of-the-art methods across various benchmarks. We implement the baseline methods based on their official implementations. The performance of the vanilla versions of LLaVA-OneVision and LLaVA-Video differs slightly from their reported results but remains within the margin of error. The best performance at the same retention ratio $R$ is highlighted in bold.
4.1 Experimental Settings
Implementation Details. We apply our method to two representative Video LLMs: LLaVA-OneVision-7B [14] is a general Video LLM, while LLaVA-Video-7B [43] employs newline tokens based on positional information of video tokens. These models encompass different architectural characteristics, ensuring a robust evaluation of our approach.
For LLaVA-OneVision, $32$ sampled frames generate a $32\times 196$ token input to the LLM, while for LLaVA-Video, $64$ sampled frames generate a $64\times 169$ token input. Our method focuses on token compression at the LLM input stage without modifying the prefilling or decoding stages. Consequently, the retention ratio $R\%$ of input tokens directly impacts inference speed and model complexity. Unless otherwise specified, we adopt the hyperparameter setting $c=8,\tau=0.9,d=0.4,p=4,\beta=0.6$ for all experiments. We conduct all evaluations using LMMs-Eval [40].
Compared Baselines. We compare FastVID with strong baselines from both image and video compression methods.
For image compression, we adopt both the widely used classic method FastV [4] and the current state-of-the-art approach VisionZip [38]. FastV prunes tokens during the prefilling stage based on attention scores from query tokens. In contrast, VisionZip prunes tokens at the input stage. However, VisionZip prunes tokens in the last layer of the vision encoder, conflicting with pooling in Video LLMs and degrading performance. To address this, we implement VisionZip, which applies pruning after pooling.
For video compression, we compare two of the latest methods, DyCoke [31] and PruneVID [11]. Both employ multi-stage pruning across the input, prefilling, and decoding stages. Although they initially prune redundant video tokens at the input stage, their prefilling and decoding processes disregard video-specific spatiotemporal redundancy. Instead, they treat video tokens as textual tokens and prune them solely based on attention scores from query tokens, similar to FastV. Furthermore, their input-stage video pruning remains inefficient, retaining $47.5\%$ of tokens in DyCoke and $50\%$ in PruneVID. This indicates that their overall efficiency is primarily influenced by prefilling and decoding stages rather than effective video redundancy removal. Since our study focuses on video spatiotemporal redundancy, we ensure a fair comparison by evaluating only their video pruning strategies.
@@ 4.1 Experimental Settings
Segmentation MVBench LongVideo- Bench MLVU VideoMME Avg.
Fixed-interval 55.1 53.6 61.7 55.4 95.3
Cluster-based 53.2 53.0 61.7 54.8 93.9
Our DySeg 55.9 56.3 62.7 57.3 98.0
Table 2: Ablation study on different video segmentation methods. As shown in Figure 3, Fixed-interval Segmentation partitions the video at equal intervals. Cluster-based Segmentation groups frames based on frame similarity. Our proposed DySeg adaptively segments the video based on transition similarity.
@@ 4.2 Comparisons with State-of-the-Art Methods
For a comprehensive evaluation, Table 1 compares our FastVID with state-of-the-art methods on benchmarks with diverse video durations. In DyCoke [31], video frames are segmented into groups of 4, retaining all tokens from the first frame while pruning the rest. Consequently, its lowest retention ratio $R$ is $25\%$. For other baselines, we conduct comparisons across different $R$ values.
Results on LLaVA-OneVision. We evaluate our FastVID against other methods on LLaVA-OneVision at retention ratios of $25\%,20\%,15\%,10\%$, comparing their token compression effectiveness. Image compression methods perform well at $R=25\%,$ but their performance declines sharply as $R$ decreases. For instance, from $R=25\%$ to $R=10\%$, FastV and VisionZip exhibit performance drops of $5.9\%$ and $9.5\%$, respectively. This indicates the limitations of spatial compression alone, which struggles to preserve essential temporal information under extreme pruning. When $R=10\%$, video compression methods, particularly PruneVID and our FastVID, significantly outperform image-based methods, demonstrating the importance of modeling temporal redundancy for efficient video understanding. While our FastVID achieves performance comparable to VisionZip at $R=25\%$, its advantage becomes increasingly evident as $R$ decreases, with an average performance drop of just $\textbf{1.0}\%$. Notably, even after pruning $\textbf{90}\%$ of the tokens, our FastVID preserves $\textbf{98.0}\%$ of the vanilla model’s performance. FastVID significantly outperforms PruneVID by $\textbf{3.1}\%$, FastV by $\textbf{7.8}\%$, and VisionZip by $\textbf{8.4}\%$, demonstrating an optimal trade-off between efficiency and accuracy.
$d$ in STPrune MVBench LongVideo- Bench MLVU VideoMME Avg.
@@
Table 3: Ablation Study on $d$ in STPrune. STPrune consists of DTM and ATS. A fraction $d$ of the pruned tokens is from DTM, while the remaining $1-d$ is from ATS. When $d=0.0$, all pruned tokens are from ATS, while at $d=1.0$, they are from DTM.
Results on LLaVA-Video. We conduct experiments on LLaVA-Video with $R=25\%$ and $R=10\%$. When $R=10\%$, the performance of both video compression methods in LLaVA-Video is lower than in LLaVA-OneVision. This discrepancy is likely due to LLaVA-Video’s higher pooling compression rate, where each frame contains $169$ tokens instead of $196$ tokens in LLaVA-OneVision. Further pruning under this higher pooling compression rate leads to a performance drop. When $R=25\%$, image compression methods achieve state-of-the-art performance on LongVideoBench. However, FastVID outperforms them significantly on other datasets, achieving superior $\textbf{98.1}\%$ performance. When $R=10\%$, FastVID consistently surpasses all other methods across all benchmarks. On LLaVA-Video, our FastVID maintains a clear advantage, outperforming other methods by $\textbf{1.4}\%$ at $R=25\%$ and $\textbf{4.0}\%$ at $R=10\%$.
4.3 Ablation Study
In this section, we conduct ablation studies on the proposed modules and hyperparameters. By default, we perform experiments on LLaVA-OneVision, setting the retention ratio $R$ to $10\%$. Under this configuration, the $32\times 196$ video tokens are compressed to an average of $19$ tokens per frame.
Ablation study on DySeg. To compare different video segmentation methods, we present qualitative and quantitative results in Figure 3 and Table 2, respectively. In Table 2, Fixed-interval Segmentation with an interval of $4$ generates $32/4=8$ segments, while Cluster-based Segmentation generates $8$ clusters, ensuring consistency with the minimal segment number $c$ in our DySeg. Fixed-interval preserves temporal order between segments, whereas Cluster-based maintains intra-segment similarity. The results show that Fixed-interval outperforms Cluster-based by $1.4\%$, highlighting the importance of temporal structure preservation in Video LLM pruning. Additionally, achieving high intra-segment similarity is also essential for efficient pruning. Our proposed DySeg successfully integrates both advantages while enabling dynamic pruning strategies, achieving a superior average performance of $\textbf{98.0}\%$.
Token Merging MVBench LongVideo- Bench MLVU VideoMME Avg.
@@ 4.3 Ablation Study
Table 4: Ablation study on different token merging strategies. Uniform, adopted in VisionZip (see Figure 1c), merges tokens through uniform sampling. Cluster-based methods (see Figure 4a) conduct clustering algorithms. Our proposed DTM employs a density-based method (see Figure 4b).
@@
Ablation study on STPrune. Our STPrune is proposed to prune tokens within each segment and consists of two key components: DTM and ATS. DTM employs density-based sampling to merge redundant tokens, preserving segment visual context. ATS leverages [CLS] attention scores to highlight important visual details. Table 3 presents an ablation study on $d$, which controls the token pruning distribution between DTM and ATS. The results show that ATS alone (when $d=0.0$) outperforms DTM alone (when $d=1.0$). However, the best performance is achieved at $d=0.4$, demonstrating that a balanced integration of both modules is crucial for preserving essential video information. This integration boosts performance by $\textbf{1.8}\%$ and $\textbf{4.5}\%$ points over ATS and DTM alone, respectively.
@@
Ablation study on DTM. In Table 4, we compare different token merging strategies for segment visual context. Our DTM consistently outperforms other approaches in both Video LLMs. Uniform struggles to distinguish objects due to indiscriminate anchor token selection (see Figure 1c). Cluster-based discards essential positional information (see Figure 4a), resulting in inferior performance, particularly in LLaVA-Video. By leveraging Density-based Sampling and Anchor-centric Aggregation, our DTM achieves superior results, especially in LLaVA-Video, where DTM outperforms other methods by $\textbf{2.9}\%$.

### v2

@@ 4 Experiments
4.1 Experimental Settings
Table 1: Comparison of state-of-the-art methods on LLaVA-OneVision [16]. The A%/Bretention ratio indicates that Aof the LLM input tokens are retained, and subsequently compressed to Bduring the LLM forward pass. The best performance among those with similar retention ratios $R$ is highlighted in bold. TFLOPs related to video tokens are reported (see the appendix for details).
Method Retention Ratio $R$ TFLOPs MVBench LongVideo Bench MLVU VideoMME Avg. Acc.
@@ 4.1 Experimental Settings
Vanilla 100% 48.82 56.9 56.4 65.2 58.6 70.3 56.6 48.8 59.3 100
DyCoke [33]${}_{\text{CVPR'25}}$ 32.5% 14.13 56.3 56.6 62.1 57.1 68.1 56.7 46.7 58.0 97.8
FastV [5]${}_{\text{ECCV'24}}$ 100%/25% 13.45 54.7 55.5 61.5 56.2 68.0 54.6 46.0 57.0 96.1
VisionZip [42]${}_{\text{CVPR'25}}$ 25% 10.73 53.7 51.2 58.5 54.1 61.6 53.4 47.2 54.4 91.7
VisionZip [42]${}_{\text{CVPR'25}}$ 25% 10.73 56.6 55.7 64.8 58.0 68.6 57.7 47.7 58.8 99.1
DyCoke [33]${}_{\text{CVPR'25}}$ 25% 10.73 49.5 48.1 55.8 51.0 61.1 48.6 43.2 51.1 86.2
FastVID 25% 10.73 56.5 56.3 64.1 58.0 69.9 56.6 47.7 58.7 99.0
FastV [5]${}_{\text{ECCV'24}}$ 100%/20% 11.38 54.1 56.6 61.2 56.2 66.8 54.6 47.2 57.0 96.1
VisionZip [42]${}_{\text{CVPR'25}}$ 19.9% 8.46 53.0 50.0 57.1 53.0 60.8 51.0 47.1 53.3 90.0
VisionZip [42]${}_{\text{CVPR'25}}$ 19.9% 8.46 55.8 55.4 64.2 58.0 68.6 57.0 48.3 58.4 98.5
FastVID 19.9% 8.46 56.3 57.1 63.9 57.9 69.3 56.7 47.7 58.8 99.1
FastV [5]${}_{\text{ECCV'24}}$ 100%/15% 9.35 53.2 54.9 59.8 54.7 65.1 53.4 45.7 55.7 93.9
VisionZip [42]${}_{\text{CVPR'25}}$ 14.8% 6.23 50.3 46.9 54.4 49.5 55.8 49.3 43.3 50.3 84.8
VisionZip [42]${}_{\text{CVPR'25}}$ 14.8% 6.23 54.3 53.9 63.1 55.5 63.0 54.4 49.1 56.7 95.6
FastVID 14.8% 6.23 56.0 56.2 63.2 57.7 69.3 56.2 47.4 58.3 98.3
FastV [5]${}_{\text{ECCV'24}}$ 100%/10% 7.36 51.7 52.1 57.7 52.4 60.9 51.4 45.0 53.5 90.2
VisionZip [42]${}_{\text{CVPR'25}}$ 9.7% 4.04 44.4 43.5 51.5 46.0 50.4 45.8 41.8 46.4 78.3
VisionZip [42]${}_{\text{CVPR'25}}$ 9.7% 4.04 51.7 48.3 59.7 52.8 59.4 52.0 46.9 53.1 89.6
PruneVID [13]${}_{\text{ACL'25}}$ 10.1% 4.23 54.2 53.8 62.3 55.9 66.4 52.9 48.3 56.6 95.4
FastVID 9.7% 4.04 55.9 56.3 62.7 57.3 67.4 56.0 48.6 58.1 98.0
PruneVID [13]${}_{\text{ACL'25}}$ 10.1%/3.9% 2.58 54.1 51.8 62.3 55.5 67.1 51.8 47.7 55.9 94.3
FastVID+FastV [5] 9.7%/3.6% 2.47 55.0 53.6 61.9 56.3 66.1 54.8 48.1 56.7 95.6
Table 2: Comparison of state-of-the-art methods on LLaVA-Video [47]. TFLOPs calculations include both video and newline tokens.
@@
Method Retention Ratio $R$ Newline Tokens $M$ TFLOPs MVBench LongVideo Bench MLVU VideoMME Avg. Acc.
Overall Short Long Score %
Vanilla 100% 832 103.2 60.4 59.6 70.3 64.1 76.9 53.4 63.6 100
DyCoke [33]${}_{\text{CVPR'25}}$ 32.1% 256 27.1 59.3 57.9 65.7 61.6 74.6 51.1 61.1 96.1
FastV [5]${}_{\text{ECCV'24}}$ 100%/25% 832/568.7 29.2 58.0 58.3 63.9 61.0 71.3 51.0 60.3 94.8
VisionZip [42]${}_{\text{CVPR'25}}$ 24.9% 64 19.5 56.4 54.1 62.1 58.6 66.3 51.2 57.8 90.9
VisionZip [42]${}_{\text{CVPR'25}}$ 24.9% 64 19.5 58.3 58.3 66.6 61.9 73.3 52.2 61.5 96.7
DyCoke [33]${}_{\text{CVPR'25}}$ 25% 208 20.7 50.8 53.0 56.9 56.1 65.8 48.9 54.2 85.2
FastVID 24.9% 64 19.5 59.3 58.3 67.7 62.6 74.9 52.0 62.0 97.5
FastVID 24.9% 715.5 24.5 59.9 57.4 68.6 63.6 74.9 53.7 62.4 98.1
FastV [5]${}_{\text{ECCV'24}}$ 100%/10% 832/311.8 16.2 55.8 55.4 58.9 57.9 67.6 48.6 57.0 89.6
VisionZip [42]${}_{\text{CVPR'25}}$ 9.5% 64 7.3 46.3 46.6 52.2 49.5 54.2 44.3 48.7 76.6
VisionZip [42]${}_{\text{CVPR'25}}$ 9.5% 64 7.3 56.6 53.6 61.7 58.7 67.6 50.1 57.7 90.7
FastVID 9.5% 64 7.3 58.3 56.2 63.9 59.6 70.9 50.7 59.5 93.6
FastVID 9.5% 508.2 10.5 58.5 56.5 64.9 60.7 71.7 51.2 60.2 94.7
Table 3: Efficiency Comparison on LLaVA-OneVision [16]. The prefilling time, defined as the latency to the first generated token, is measured on VideoMME using an A100 GPU.
Method # Token TFLOPs Prefilling Time (ms) Avg. Acc.
Segmentation Compression LLM Forward Total
Vanilla 6272 (100%) 48.82 (100%) 476.3 476.3 (1.0$\times$) 59.3 (100%)
PruneVID [13] 635.6 (10.1%) 4.23 (8.7%) 5.2 32.0 64.3 101.5 (4.7$\times$) 56.6 (95.4%)
FastVID 608 (9.7%) 4.04 (8.3%) 0.5 5.6 61.1 67.2 (7.1$\times$) 58.1 (98.0%)
Implementation Details. We apply our method to two representative Video LLMs: LLaVA-OneVision-7B [16] is a general Video LLM, while LLaVA-Video-7B [47] employs newline tokens based on positional information of video tokens. These models encompass different architectural characteristics, ensuring a robust evaluation of our approach. Unless otherwise specified, we adopt the hyperparameter setting $c=8,\tau=0.9,d=0.4,p=4,\alpha=0.6$ for all experiments. For LLaVA-OneVision, $32$ sampled frames generate a $32\times 196$ token input to the LLM. We test with $r\in\{25\%,20\%,15\%,10\%\}$. For LLaVA-Video, $64$ sampled frames generate a $64\times 169$ token input. We test with $r\in\{25\%,10\%\}$. We conduct all evaluations using LMMs-Eval [44] on A100 GPUs.
Compared Baselines. (1) For image compression, we adopt both the widely used classic method FastV [5] and the current state-of-the-art approach VisionZip [42]. VisionZip prunes tokens in the last layer of the vision encoder, conflicting with pooling in Video LLMs and degrading performance. To address this, we implement VisionZip, which applies pruning after pooling. (2) For video compression, we compare two recent methods, DyCoke [33] and PruneVID [13]. Our FastVID focuses on pruning during the prefilling stage. To ensure fairness, we reimplement these baselines without pruning in the decoding stage. PruneVID applies two-stage pruning on both input tokens and the LLM’s 10th layer. PruneVID refers to the variant that prunes only input tokens. More implementation details of these baselines are provided in the appendix.
@@ 4.2 Comparisons with State-of-the-Art Methods
For a comprehensive evaluation, we compare our FastVID with state-of-the-art methods on benchmarks with diverse video durations. In DyCoke, video frames are evenly divided into 4-frame segments, retaining all tokens from the first frame while pruning the rest. Consequently, its lowest retention ratio $R$ is $25\%$. For other baselines, we conduct comparisons across different $R$ values.
Results on LLaVA-OneVision. In Table 1, we evaluate our FastVID against other methods on LLaVA-OneVision. While FastV and VisionZip perform well at $R=25\%$, but their performance declines sharply as $R$ decreases. This indicates the limitations of spatial compression alone, which struggles to preserve essential temporal information under extreme pruning. Notably, even after pruning 90.3% of the tokens, our FastVID preserves 98.0% of the vanilla model’s performance. To compare with PrunVID, we additionally apply FastV at the 10th layer of the LLM. FastVID+FastV achieves 95.6% of the original accuracy at 2.47 (5.1%) TFLOPs.
Table 4: Ablation Study on $d$ in STPrune. STPrune consists of DTM and ATS. A fraction $d$ of the retained tokens is from DTM, while the remaining $1-d$ is from ATS.
$d$ in STPrune MVBench LongVideo- Bench MLVU VideoMME Avg. Acc.
@@
Table 5: Ablation study on different token merging strategies. Figure 1(b1) presents Uniform. Figure 4(a) presents Cluster-based.
Token Merging MVBench LongVideo- Bench MLVU VideoMME Avg. Acc.
@@
Results on LLaVA-Video. LLaVA-Video adopts a unique design by inserting newline tokens after each height-wise position in every frame, producing $64\times 13\times(13+1)$ tokens. In Table 2, for all methods, when the positional information of retained tokens is preserved, we also retain the associated newline tokens. For a fair comparison with VisionZip, we evaluate FastVID, where only one newline token is retained per frame. Notably, our FastVID consistently outperforms all baselines across various retention ratios.
[…]
4.3 Ablation Study
By default, we conduct ablation studies on LLaVA-OneVision at $r=10\%$. Further hyperparameter analysis is provided in the appendix.
Table 6: Ablation study on video segmentation.
Segmentation MVBench LongVideo- Bench MLVU VideoMME Avg. Acc.
Fixed-interval 55.1 53.6 61.7 55.4 95.3
Cluster-based 53.2 53.0 61.7 54.8 93.9
Our DySeg 55.9 56.3 62.7 57.3 98.0
Ablation study on DySeg. To compare different video segmentation methods, we present qualitative and quantitative results in Figure 3 and Table 6, respectively. In Table 6, Fixed-interval Segmentation with an interval of $4$ generates $32/4=8$ segments, while Cluster-based Segmentation generates $8$ clusters, ensuring consistency with the minimal segment number $c$ in our DySeg. Fixed-interval preserves temporal order between segments, whereas Cluster-based maintains intra-segment similarity. The results show that Fixed-interval outperforms Cluster-based by $1.4\%$, highlighting the importance of temporal structure preservation in Video LLM pruning. Additionally, high intra-segment similarity is also essential for effective pruning. Our proposed DySeg successfully integrates both advantages while enabling dynamic pruning, achieving a superior average performance of 98.0%.
@@ 4.3 Ablation Study
Ablation study on STPrune. STPrune is proposed to prune tokens within each segment and consists of two key components: DTM and ATS. DTM employs density-based sampling to merge redundant tokens, preserving segment visual context. ATS leverages [CLS] attention scores to highlight important visual details. Table 5 presents an ablation study on $d$, which controls the pruning distribution between DTM and ATS. The results show that ATS alone (when $d=0.0$) outperforms DTM alone (when $d=1.0$). However, the best performance is achieved at $d=0.4$, demonstrating that a balanced integration of both modules is crucial for preserving essential video information. This integration boosts performance by 1.8% and 4.5% points over ATS and DTM alone, respectively.
@@
Ablation study on DTM. In Table 5, we compare different token merging strategies for segment visual context. Uniform, used in VisionZip, selects anchor tokens in a content-agnostic manner, which limits its ability to distinguish semantically meaningful objects. Cluster-based discards essential positional information, resulting in inferior performance, particularly in LLaVA-Video. By leveraging Density-based Sampling and Anchor-centric Aggregation, our DTM achieves superior results, especially in LLaVA-Video, where DTM outperforms other methods by 2.9%.
Figure 5: Visualization of FastVID with $c=0,\tau=0.9,d=0.4,p=2$. We retain a total of 40 tokens, including 16 tokens in DTM (highlighted in red) and 24 in ATS (highlighted in blue). In DTM, patches that share the same inner and border color are merged.
4.4 Qualitative Results
Figure 5 visualizes the proposed DySeg and STPrune. In Figure 5(a), all frames are grouped into a single segment, with the 1st and 3rd frames selected as anchor frames, where DTM is applied. In Figure 5(b), the frames are divided into two segments, with the 1st, 3rd, and 4th frames selected as anchor frames. Blue tokens represent those selected by ATS due to high [CLS] attention. These tokens readily cluster together and do not always correspond to object regions, suggesting that while they reflect salient [CLS] information, they lack broader visual context. In contrast, DTM’s red tokens effectively aggregate visually similar content across frames, significantly reducing spatiotemporal redundancy while preserving visual context. Together, ATS and DTM offer complementary benefits for effective video compression. More qualitative results are provided in the appendix.
5 Conclusion
In this paper, we introduced FastVID, a novel inference-time pruning framework designed to accelerate Video LLMs by effectively reducing spatiotemporal redundancy. Through a comprehensive analysis of video tokens from both temporal context and visual context, FastVID dynamically partitions videos into temporally ordered segments and employs density-based token pruning within each segment. Extensive experiments across multiple Video LLMs and benchmarks demonstrate its generalization ability and effectiveness. Crucially, FastVID maintains high performance even under extreme compression rates, enabling practical deployment of fast Video LLMs.
Appendix
This appendix provides additional details and results to support our main paper:
  • Section A presents additional experimental settings, including reproduction details of the compared baselines and an estimation of the computational cost.
  • Section B presents additional experimental results, such as ablation studies on key hyperparameters, further qualitative results, and results on Qwen2-VL [37].
  • Section C discusses limitations of our approach and its potential broader impacts.
Appendix A Additional Experimental Settings
A.1 Reproduction Details of Compared Baselines
All experiments are conducted using LMMs-Eval [^2] https://github.com/EvolvingLMMs-Lab/lmms-eval, MIT License [44] for consistency. The performance of the vanilla versions of LLaVA-OneVision [^3] https://github.com/LLaVA-VL/LLaVA-NeXT, Apache License 2.0 [16], LLaVA-Video33footnotemark: 3 [47], and Qwen2-VL [^4] https://github.com/QwenLM/Qwen2.5-VL, Apache License 2.0 [37] differs slightly from their reported results, remaining within an acceptable margin of error. We reimplement all baseline methods using LMMs-Eval, following their official implementations:
  • FastV [^5] https://github.com/pkunlp-icler/FastV [5] (ECCV 2024). FastV performs token pruning at the $K$-th layer of the LLM using attention scores, with a filtering ratio $R$. We reimplement it with $K=2$, using $R\in\{75\%,80\%,85\%,90\%\}$ in Table 1 and $R\in\{75\%,90\%\}$ in Table 2.
  • VisionZip [^6] https://github.com/dvlab-research/VisionZip, Apache License 2.0 [42] (CVPR 2025). Visionzip performs pruning at the vision encoder’s output, which conflicts with pooling operations in Video LLMs and degrades performance. To address this, we implement VisionZip, which applies pruning after pooling. Following the original settings, each frame retains both dominant and contextual tokens in a $54\!:\!10$ ratio. We define $r$ as the proportion of tokens retained per frame. We set $r\in\{25\%,20\%,15\%,10\%\}$ in Table 1 and $r\in\{25\%,10\%\}$ in Table 2.
  • DyCoke [^7] https://github.com/KD-TAO/DyCoke, Apache License 2.0 [33] (CVPR 2025). DyCoke prunes in both prefilling and decoding stages. For fair comparison, we only evaluate its pre-filling stage. The pruning rate in the TTM module is set $K\in\{0.9,1.0\}$ in both Table 1 and Table 2.
  • PruneVID [^8] https://github.com/Visual-AI/PruneVid, CC BY-NC-SA 4.0 License [13] (ACL 2025). PruneVID includes both input-stage and intra-LLM pruning in the pre-filling phase, along with decoding pruning. For fair comparison, we evaluate two variants: PruneVID (input-stage pruning only) and PruneVID (full prefilling pruning). Following the original settings, we use a threshold $\tau=0.8$, temporal segment ratio $\gamma=0.25$, token selection ratio $\alpha=0.4$, and attention calculations use the 10th layer. The cluster ratio $\beta$ controls compression in the input-stage pruning. We use $\beta=11\%$ in Table 1 and Table 3.
A.2 Computational Cost Estimation
Following prior works [5, 39], we report the theoretical FLOPs of the LLM related to visual/video tokens. Specifically, LLaVA-OneVision [16], LLaVA-Video [47], and Qwen2-VL [37] are all built on Qwen2 [41], which employs grouped-query attention [2] and a three-layer FFN structure. The per-layer FLOPs of the LLM are computed as:
$2nD(h_{kv}d)+2nD^{2}+2n^{2}D+3nDD^{\prime}$ (6)
where $n$ is the number of video tokens, $D$ is the hidden state size, $D^{\prime}$ is the FFN intermediate size, $h_{kv}$ is the number of key/value heads, and $d$ is the head dimension.
Appendix B Additional Experimental Results
B.1 Ablation study on $c$ and $\tau$ in DySeg